OpenStudy (anonymous):

I GIVE MEADALS...PLEASE HELP How many solutions does the equation 2|x − 3| + 4 = 6 have?

OpenStudy (anonymous):

First subtract 4 from both sides, and then divide by 2, you get /x-3/ = 1 so x- 3 = 1 and x-3 = -1 x = 4 and x = 2 TWO solutions.

OpenStudy (anonymous):

options: A.) zero B.) one C.) two D.) Infinitely Many

OpenStudy (anonymous):

See my solutions above.

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

welcome.

OpenStudy (anonymous):

i hv another 1...ill post over to the left